Opened 12 years ago

Last modified 9 years ago

#364 assigned flight processing

EU10/03, flight day 234c/2010, Lake Balaton East Area A — at Version 12

Reported by: iopa Owned by: knpa
Priority: alpha 5 Milestone: 2010 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by benj)

Data location: ~arsf/workspace/EU10_03-2010_234c_Lake_Balaton_East_Area_A

Data arrived from ARSF via SATA disk Q on 15/09/2010.

Scientific objective: Lake Balaton provides an ideal area for testing emerging algorithms on retrieval of in-water constituents and on mapping bloom-forming cyanobacteria using in-situ radiometric, airborne and satellite data.

Priority: a5

PI: Andrew Tyler

Note: Data is currently symlinked from ~arsf/arsf_data/2010/flight_data/archive_overflow and when this is resolved the symlinks will need to be re-directed.

Sensors:

  • Eagle (27/10/2010)
  • Hawk (Requested by PI)
  • Leica LIDAR (Requested but low priority)
  • Digital Photography (27/10/2010)

Change History (12)

comment:1 Changed 12 years ago by iopa

  • Description modified (diff)

comment:2 Changed 12 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

Started eagle processing.

comment:3 Changed 12 years ago by iopa

Applanix and IPAS have been done. The applanix data currently in the workspace covers both this project and also 234b.
For IPAS, there is processed IPAS in 234a in the workspace that covers 234a,234b,234c.
The proc directory from there could be entirely copied over to 234c or just symlinked, your choice.

comment:4 Changed 12 years ago by iopa

Taken camera processing.

comment:5 Changed 12 years ago by iopa

Camera ready for checking.

All images were processed, the test images were taken earlier in the day.

comment:6 Changed 12 years ago by emca

Delivery checked camera imagery:
~airborne/workspace/EU10_03-2010_234c_Lake_Balaton_East_Area_A/delivery/20101008/EU10_03

All data looks good. Ready to deliver.

comment:7 Changed 12 years ago by knpa

Found Eagle and Hawk SCTs:

eagle hawk
1 0.03 -0.01
2 0.05 0.00
3 0.00 0.00
4 0.04 0.00
5 0.00 -0.01
6 -0.01 -0.02

comment:8 Changed 12 years ago by knpa

Appears to be an issue with the bandset labels in the header files; they start at 900nm which is lower than it should be. Possibly incorrect bandset selected or none specified. Will leave this until end of the year and new Hawk calibration.

comment:9 Changed 12 years ago by knpa

Eagle awaiting delivery check.

comment:10 Changed 12 years ago by iopa

Eagle has been delivery checked;
the only issue found was that fastQC gives a header-filesize error when opening e234c061.bil
but everything seems to be in order between the header values and the filesize.

Moved .tex and .cfg files out of the delivery directory.

Eagle is ready to deliver.

comment:11 Changed 12 years ago by emca

  • Description modified (diff)

Eagle and Camera dispatched to PI 27/10/2010

comment:12 Changed 12 years ago by benj

  • Description modified (diff)
Note: See TracTickets for help on using tickets.